Determine which of the polynomials below is (are) irreducible over Q. a. x5+9x4+12x2+6 b. x4+x+1


1
Expert's answer
2021-11-26T05:10:17-0500

Solution:

(a) We can check whether "x^5+9x^4+12x^2+6" is reducible over Q or not by Eisenstein's criterion.

Let "p=3" (a prime number). We know that "a_n" is the coefficient of highest power. Now, we notice that "3 \nmid a_n=1", but since "3|9, 3|12, 3|6" and "3|0", i.e., "3" divides all the other coefficients of given polynomial.

We also notice that "3^2=9 \nmid 6=a_0"

Thus, Eisenstein's criterion is satisfied and so, "x^5+9x^4+12x^2+6" is irreducible over Q[x].

Answer"x^5+9x^4+12x^2+6" is irreducible over Q[x].


(b)We apply the mod 2 test.

"x^4+x+1\in Z[x]" is primitive. Thus "x^4+x+1\in Q[x]" is irreducible if and only if "x^4+x+1\in Z[x]" is irreducible.

The mod 2 reduction of "x^4+x+1\in Z[x]" is "f(x)=x^4+x+1\in Z_2[x]" . Since

"f(a)=1 \neq0" for all "a\in Z2~" it follows that "f(x)" has no linear factors.

Suppose that "f(x)" is reducible. Then it must be the product of quadratic factors. There are 3 quadratic reducible polynomials in "Z_2[x]".Therefore there is 1 irreducible quadratic in "Z_2[x]" which is "x^2 + x + 1" since this polynomial has no roots in "Z_2[x]". Therefore "f(x)=(x^2+x+1)^2=x^4+x^2+1" which is not the case.

We have shown "f(x)\in Z_2[x]" is irreducible. Thus "x^4+x+1\in Z[x]" is irreducible

which means "x^4+x+1\in Q[x]" is also.

Answer: "x^4+x+1" irreducible over Q[x].